Litres: Beginning work in process inventory 2,800 litres Started production 4,620 litres Completed and transferred out...

Litres:
Beginning work in process inventory 2,800 litres
Started production 4,620 litres
Completed and transferred out to Packaging in March 6,420 litres
Ending work in process inventory (80% of the way through
the fermenting process) 1,000 litres
Costs:
Beginning work in process inventory ($2,200 of direct
materials and $1,462 of conversion cost) $3,662
Costs added during March:
Direct materials 10,414
Direct labour 950
Manufacturing overhead 1,920
Total costs added during March $13,284

1.

Fill in the time line for the Fermenting Department.

2.

Summarize the flow of physical units and compute the total equivalent units.

3.

Summarize total costs to account for and compute the cost per equivalent unit for direct materials and conversion costs.

4.

Assign total costs to units​ (litres):

a. Completed and transferred out to the Packaging Department

b. In the Fermenting Department ending work in process inventory

5.

What is the average cost per litre transferred out of Fermenting into​ packaging? Why would JohnsonJohnson ​Vineyard's managers want to know this​ cost?

Requirement 1. Fill in the time line for the Fermenting Department.

Start

100%

Complete

Complete

Transferred out to Packaging

Requirement 2. Summarize the flow of physical units and compute the total equivalent units.

Johnson Vineyards

Fermenting Department

Flow of Physical Units and Computation of Equivalent Units

Flow of

Equivalent Units

Physical

Direct

Conversion

Flow of Production

Units

Materials

Costs

Units to account for:

Units accounted for:

Total equivalent units

Requirement 3. Summarize total costs to account for and compute the cost per equivalent unit for direct materials and conversion costs. ​(Round the cost per equivalent unit to the nearest​ cent.)

Johnson Vineyards

Fermenting Department

Cost per Equivalent Unit

Direct

Conversion

Materials

Costs

Total

Cost per equivalent unit

Requirement 4. Assign total costs to units​ (litres):

Now we can calculate the total cost of the units for a. completed and transferred out to the Packaging Department and b. the Fermenting Department ending work in process. ​(Enter quantities​ first, then the cost per equivalent unit amounts in the same order as calculated in the preceding step. Round your answers to the nearest​ cent.)

Johnson Vineyards

Fermenting Department

Assignment of Costs

Assign costs:

Direct Materials

Conversion Costs

Total

a.

x (

+

)

b.

x

x

Total cost accounted for

Requirement 5. What is the average cost per litre transferred out of Fermenting into​ Packaging? Why would Johnson ​Vineyard's managers want to know this​ cost?

First identify the labels to compute the average cost per litre transferred​ out, then compute the average cost per unit. ​(Round your answer to the nearest​ cent.)

Avg. cost

/

=

per litre

/

=

Why would Johnson ​Vineyard's managers want to know this​ cost?

Managers will want to know this cost to compare it to their .................budgeted target costs cost per equivalent unit direct material costs . They may also use the cost information when ..................

.
Answer #1
1to 4) Production report of Fermenting Department for the month March:
(no % completion is provided, so weighted average costing method is adopted)
Phy units Material Conversion costs Total
Open WIP 2800 2800 2800
Started and completed 3620 3620 3620
Transferred out to packing 6420
Close WIP 1000 1000 800 (80%)
Equivalent units 7420 7220
Total Costs:
Open WIP 2200 1462
Incurred during month 10414 2870
Total Costs to Acct for: 12614 4332 16946
Cost per Eq unit 1.7 0.6
Cost of Transferred out 6420*1.70=10914 6420*0.60=3852 14766
Cost of Close WIP 1000*1.7=1700 800*0.60=480 2180  
Total Costs acctd for: 16946

5) Average cost per lit. transferred out = Total cost transferred out/Total quantity transferred out

= 14766/6420 = $2.30/liter

Managers will want to know this cost to compare it to their Production budgeted target costs cost per equivalent unit direct material costs . They may also use the cost information when to know the material cost comparison between two periods.